### v1

@@ 4.1.1 CrisisMMD
The CrisisMMD dataset [34] is a valuable and comprehensive resource in the field of disaster management. It consists of a vast collection of multimodal social media posts that encompass texts and images from seven natural disaster events. The CrisisMMD dataset serves as a benchmark for evaluating the performance of state-of-the-art methods in disaster management. The number of samples from the CrisisMMD dataset utilized in our study is displayed in Table 1.
@@ 4.1.2 TSEqD
Due to the dearth of multimodal datasets in the field of disaster management. We have curated a vast collection of social media posts comprising of texts and images from the Turkey-Syria Earthquake. The dataset, TSEqD (Turkey-Syria Earthquake Dataset), has two different sets of annotations: one for Informative Tasks and another for Humanitarian Tasks. We have chosen to adhere to the CrisisMMD dataset’s consistent labeling format. The process of data curation, annotation, and the distribution of labels across various classes have been elaborated in subsequent sections.
@@ 4.1.2.1 Data Curation:
4.1.2.2 Multitask Annotation:
We first discuss the annotation process for Task 1 and then Task 2.
4.1.2.2.1 Annotation Process for Task 1 (Informative Task):
@@ 4.1.2.2.2 Annotation Process for Task 2 (Humanitarian Task):
The distribution of text and image samples in each class in TSEqD dataset is shown in Table 4 In brief, we have annotated the dataset for various tasks, such as the Informative Task (Task 1) and Humanitarian Task (Task 2). The annotated dataset, namely TSEqD, contains about 10,352 data samples. The dataset contains 22 columns and contains necessary meta information which can be helpful to the researchers. To the best of our knowledge, no such dataset in disaster management has this unique and meaningful data.
@@
4.1.2.3 Inter-Annotator Agreement:
In the annotation process, we conducted separate annotations for both image and text data. The inter-annotator agreement scores for different modalities under different tasks can be seen in Fig. 3. Task 1, being a binary task, displays a notably higher inter-annotator agreement, indicating minimal ambiguity in the annotation process. In contrast, Task 2 exhibits a substantial decrease in agreement, primarily due to two significant factors. First, Task 2 is dependent on Task 1, and second, the inherent complexity of Task 2 leads to various possible interpretations. Notably, Task 2 may involve cases where a single image corresponds to multiple labels, potentially causing discrepancies between annotators. While both tasks exhibit a moderate level of agreement, Task 2, which focuses on humanitarian images, shows significantly lower agreement compared to humanitarian text. The level of agreement and the reliable data percentage after the annotation process are shown in the Table 5. The inter-annotator agreement is quantified using Cohen’s Kappa coefficient, a statistical measure that accounts for chance agreement, and it is computed as follows.
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (29)
As shown in Eq. 29, $P_{o}$ is the actual agreement of the annotators. It is calculated by comparing the number of times annotators had the same annotations for given data.
@@ 4.2 Creation of Crisis Lexicons
In the domain of crisis management, a well-constructed crisis lexicon list is essential, facilitating the clear and accurate communication of vital information. By integrating crisis-specific keywords in social media posts, stakeholders can efficiently sift through extensive online content, accessing pertinent information directly linked to the ongoing crisis. Olteanu et al. [50] formulated a lexicon list tailored to floods, encompassing 360 crisis-related terms. Additionally, we sourced 400 crisis lexicons from various online sources, such as Wikipedia [^4] https://en.wikipedia.org/wiki/Earthquake, and Disaster Words Vocabulary-List [^5] https://www.vocabulary.com/lists/32961. Our innovative approach involved leveraging a text encoder to generate embeddings for these meticulously curated crisis lexicons, along with word-level embeddings for tweets. By computing the cosine similarity between the crisis lexicon embeddings and tweet embeddings, we pinpointed words surpassing a threshold of 0.8, subsequently integrating them into the crisis lexicon list. These comprehensive crisis lexicon lists offer domain-specific insights, potentially enhancing the ability to discern pertinent content shared on social media platforms.
@@ 4.3 Comparison Methods
BERT: BERT (Bidirectional Encoder Representations from Transformers) [37] model is pre-trained on vast corpora, enabling it to capture contextual nuances and semantics, making it exceptionally effective for vast NLP tasks, including question-answering, text classification, and language understanding.
@@
ResNet: ResNet (Residual Network) [51] model addresses the issues of vanishing gradient in deep convolutional neural networks by introducing residual connections. ResNet has established itself as a cornerstone in the field of computer vision by achieving state-of-the-art performance in image recognition tasks.
@@
VisualBERT: VisualBERT [52] is a groundbreaking model that combines vision and language understanding by integrating the Transformer architecture with both convolutional and recurrent neural networks. It leverages pre-trained BERT embeddings alongside visual features, enabling multimodal tasks such as visual question answering and image captioning.
@@
VilBERT: Vision and Language BERT (VilBERT) [53] model is a pioneering architecture that integrates visual and textual information within a single BERT-based framework. Through pre-training on vast amounts of multimodal data, ViLBERT captures nuanced relationships between vision and language, enabling it to excel in tasks like visual question answering, image-text retrieval, and image captioning.
@@
VilT: The VilT (Vision-and-Language Transformer) [54] model harnesses the power of transformer architecture to seamlessly fuse textual and visual information. By jointly pre-trained on different vision and language data, VilT achieves remarkable capabilities in tasks like image captioning, visual question answering, and cross-modal retrieval, offering a versatile and powerful solution for bridging the gap between vision and language domains.
@@
SCBD: SCBD (SSE-Cross-BERT-DenseNet) [18] model presents an approach to real-time event detection, particularly in emergency response scenarios. The model introduces a cross-attention module to filter out irrelevant components from text and image inputs, addressing the limitations of unimodal approaches. Additionally, SCBD employs a multimodal graph-based approach, allowing stochastic transitions between embeddings of different multimodal pairs during training to improve regularization and handle limited training data effectively.
@@
MANN: Multimodal Adversarial Neural Network (MANN) [19] addresses a significant challenge in disaster-related message detection on social media by introducing an approach to identifying posts related to emerging disaster events. It has the ability to extract comprehensive features from both text and images, enabling accurate detection of disaster-related content and adaptability to unforeseen events through adversarial training.
@@
CAMM: The Cross-Attention Multi-Modal (CAMM) [55] deep neural network is proposed for classifying multimodal disaster data on social media. Integrating text and image information from Twitter posts using attention mechanisms, CAMM employs the attention mask from the textual modality to highlight relevant features in the visual modality. This cross-attention-based method emphasizes its efficacy in capturing pertinent cross-domain features in both text and image data.
@@
RoBERTaMFT: RoBERTa-based multimodal fusion transformer [57], addresses the challenges of multimodal learning. Its co-learning methodology effectively addresses issues related to recognition performance limitations and data imbalances among modalities. RoBERTaMFT is poised to offer significant advancements in multimodal understanding and classification by promoting balanced and effective learning across different data modalities.
DMCC: Deep Multimodal Crisis Categorization Framework (DMCC) [58] employs a two-level fusion strategy to seamlessly integrate textual and visual data. At the feature level, the MCA block ensures an effective fusion of these modalities. This holds the potential to enhance real-time crisis assessment and resource allocation, offering valuable support to emergency responders in dynamic and evolving situations.
@@ 4.5.1 Experimental Setup
In this study, all neural models are implemented using the PyTorch and TensorFlow frameworks, executed on a Linux server equipped with an Intel(R) Xeon(R) Silver 4215R CPU @3.20 GHz, 256GB of RAM, and a 16GB NVIDIA Tesla T4 GPU. The pre-trained BERT, CLIP, and ResNet models are accessible through the Hugging Face repository. We adopt the Adam optimizer and set the learning rate hyper-parameter for the model to $55e-$ 5. For the Feature Extraction and Interaction module (FEIM), we configure the dimensionality of hidden representations to $d=768$ for the textual modality and $1024$ for the visual modality. Similarly, for the Multimodal Graph Learning module (MGLM), the dimensionality of hidden representations for both textual and visual modalities is set to $d=512$. To mitigate overfitting, a dropout rate of $0.2$ is applied. The training process is conducted over 100 epochs.
4.5.2 Network Hyperparameters
@@
In this section, we systematically explore hyperparameters for our proposed model, such as learning rate, epochs, dropout, batch size, and optimizer, with the goal of identifying optimal configurations for improved training and overall performance. For the learning rate, we conducted comprehensive experiments across a range from $15e-$ 1 to $75e-$ 7 to identify the most effective rate for optimal model performance. Subsequently, after thorough evaluation, we determined that a learning rate of $55e-$ 5 yielded the best results, and this value was consequently adopted as part of our finalized configuration. The exploration of epochs involved intentional variations, considering values of 50, 80, 100, and 120. Through systematic experimentation, we observed that setting the number of epochs to 100 provided the optimal balance between model convergence and performance enhancement. As a result, 100 epochs were selected and incorporated into our final hyperparameter configuration. In addressing the batch size, we systematically evaluated sizes of 16, 32, 64, and 128 to accommodate diverse computational resources and memory constraints. After a thorough assessment, a batch size of 32 emerged as the choice that delivered the best results, aligning with our commitment to optimizing model performance under varied conditions. The exploration of dropout rates involved testing values within the range of $0.2$ to $0.5$. After careful consideration of performance metrics, a dropout rate of $0.2$ was identified as optimal, providing an effective regularization mechanism without compromising the model’s learning capacity. Throughout the experimentation process, the optimizer emerged as a critical component influencing model performance. After evaluating various options, Adam stood out as the preferred optimizer, demonstrating significant performance in optimizing the learning process of our model.
In summary, our finalized hyperparameter configuration includes a learning rate of $55e-$ 5, a selected batch size of 32, 100 epochs, a dropout rate of 0.2, and the utilization of the Adam optimizer during the model training process. These choices represent the culmination of rigorous experimentation aimed at achieving the optimal balance in model performance.
@@ 4.5.3 Performance Comparison on CrisisMMD
To determine how well our proposed model performs, we conducted a comprehensive evaluation by comparing CrisisSpot to both unimodal and multimodal state-of-the-art methods. We used the CrisisMMD dataset [34] for this assessment. We measured the performance of these methods using well-established metrics as discussed in subsubsection 4.5.3.The results present in Table 6 and Table 7 clearly show that CrisisSpot significantly outperforms both the unimodal and multimodal state-of-the-art methods. These findings provide strong evidence to support the effectiveness of our proposed approach.
4.5.3.1 Performance Comparison for Informative Task
To affirm the effectiveness of CrisisSpot in addressing the Informative Task on the CrisisMMD dataset [34], we carried out a thorough comparative analysis against unimodal and multimodal state-of-the-art approaches, as outlined in Table 6. Our proposed approach yielded impressive results, with an accuracy of 97.58% and an F1-score of 98.23% for the Informative Task. Notably, our approach surpassed both the unimodal and multimodal state-of-the-art benchmarks by a significant margin.
Transformer-based models, such as VisualBERT [52], ViLT [54], and ViLBERT [53], are renowned for their capabilities in addressing multimodal classification problems by efficiently handling text-image pairs and providing enhanced contextual embeddings. In our proposed approach, we harness the power of these transformer-based techniques, coupled with the incorporation of text-semantic scores and user engagement metrics within our model. Notably, our method surpasses VisualBERT by 9.48%, 10.41%, 12.57%, and 11.47% across all evaluation metrics. Similarly, our method, CrisisSpot, outperformed ViLBERT, achieving an accuracy improvement of 9.33%, precision enhancement of 7.78%, recall improvement of 9.66%, and an F1-score increase of 8.70%. Furthermore, in comparison to ViLT, our approach excelled with an accuracy improvement of 9.98%, precision improvement of 8.41%, recall enhancement of 10.62%, and an F1-score increase of 8.50%.
@@ 4.5.3.1 Performance Comparison for Informative Task
Table 6: Performance Comparison on CrisisMMD for Informative Task
Model Acc Prec Rec F1
BERT [37] 86.22 86.01 86.22 86.11
ResNet [51] 84.88 84.76 84.89 84.82
Visual BERT [52] 88.10 86.29 87.23 86.76
VilT [54] 87.60 88.29 89.18 88.73
ViLBERT [53] 88.25 88.92 90.14 89.53
SCBD [18] 88.52 91.13 91.84 91.48
MANN [19] 84.09 86.38 90.58 88.43
CAMM [55] 87.22 89.56 91.65 90.59

### v2

@@ 5.1.1 CrisisMMD
The CrisisMMD dataset (Alam et al., 2018b) is a valuable and comprehensive resource in the field of disaster management. It consists of a vast collection of multimodal social media posts that encompass texts and images from seven natural disaster events. The CrisisMMD dataset serves as a benchmark for evaluating the performance of state-of-the-art methods in disaster management. The number of samples from the CrisisMMD dataset utilized in our study is displayed in Table 1.
@@ 5.1.2 TSEqD
Due to the dearth of multimodal datasets in the field of disaster management. We have curated a vast collection of social media posts comprising of texts and images from the Turkey-Syria Earthquake [^4] https://github.com/Shahid-135/CrisisSpot. The dataset, TSEqD (Turkey-Syria Earthquake Dataset), has two different sets of annotations: one for Informative Tasks and another for Humanitarian Tasks. We have chosen to adhere to the CrisisMMD dataset’s consistent labeling format. The process of data curation, annotation, and the distribution of labels across various classes have been elaborated in subsequent sections.
@@ 5.1.2.1 Data Curation:
5.1.2.2 Multitask Annotation:
We will begin by examining the annotation process for the Informative task (Task 1), which involves two labels: informative and non-informative. Following that, we will address the Humanitarian task (Task 2), which encompasses eight labels as discussed in subsequent sections.
5.1.2.2.1 Annotation Process for Task 1 (Informative Task):
@@ 5.1.2.2.2 Annotation Process for Task 2 (Humanitarian Task):
The distribution of text and image samples in each class in the TSEqD dataset is shown in Table 4 In brief, we have annotated the dataset for various tasks, such as the Informative Task (Task 1) and Humanitarian Task (Task 2). The annotated dataset, namely TSEqD, contains about 10,352 data samples. The dataset contains 22 columns and contains necessary meta information that can be helpful to the researchers. To the best of our knowledge, no such dataset in disaster management has this unique and meaningful data.
5.1.2.3 Inter-Annotator Agreement:
In the annotation process, we conducted separate annotations for both image and text data. The inter-annotator agreement scores for different modalities under different tasks can be seen in Fig. 3. Task 1, being a binary task, displays a notably higher inter-annotator agreement, indicating minimal ambiguity in the annotation process. In contrast, Task 2 exhibits a substantial decrease in agreement, primarily due to two significant factors.
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (30)
@@ 5.1.2.3 Inter-Annotator Agreement:
First, Task 2 is dependent on Task 1, and second, the inherent complexity of Task 2 leads to various possible interpretations. Notably, Task 2 may involve cases where a single image corresponds to multiple labels, potentially causing discrepancies between annotators. While both tasks exhibit a moderate level of agreement, Task 2, which focuses on humanitarian images, shows significantly lower agreement compared to humanitarian text. The level of agreement and the reliable data percentage after the annotation process are shown in the Table 5.
@@
The inter-annotator agreement is quantified using Cohen’s Kappa coefficient, a statistical measure that accounts for chance agreement, and it is computed as follows. As shown in Eq.˜30, $P_{o}$ is the actual agreement of the annotators. It is calculated by comparing the number of times annotators had the same annotations for given data.
@@ 5.2 Creation of Crisis Lexicons
In the domain of crisis management, a well-constructed crisis lexicon list is essential, facilitating the clear and accurate communication of vital information. By integrating crisis-specific keywords in social media posts, stakeholders can efficiently sift through extensive online content, accessing pertinent information directly linked to the ongoing crisis. Olteanu et al. (Olteanu et al., 2014) formulated a lexicon list tailored to floods, encompassing 360 crisis-related terms. Additionally, we sourced 400 crisis lexicons from various online sources, such as Wikipedia [^5] https://en.wikipedia.org/wiki/Earthquake, and Disaster Words Vocabulary-List [^6] https://www.vocabulary.com/lists/32961. Our innovative approach involved leveraging a text encoder to generate embeddings for these meticulously curated crisis lexicons, along with word-level embeddings for tweets. By computing the cosine similarity between the crisis lexicon embeddings and tweet embeddings, we pinpointed words surpassing a threshold of 0.8, subsequently integrating them into the crisis lexicon list. These comprehensive crisis lexicon lists offer domain-specific insights, potentially enhancing the ability to discern pertinent content shared on social media platforms.
@@ 5.3 Comparison Methods
BERT: BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2019) model is pre-trained on vast corpora, enabling it to capture contextual nuances and semantics, making it effective for vast NLP tasks, including question-answering, text classification, and language understanding.
@@
ResNet: ResNet (Residual Network) (He et al., 2016) model addresses the issues of vanishing gradient in deep convolutional neural networks by introducing residual connections. ResNet has established itself as a cornerstone in the field of computer vision by achieving state-of-the-art performance in image recognition tasks.
@@
VisualBERT: VisualBERT (Li et al., 2019) is a groundbreaking model that combines vision and language understanding by integrating the Transformer architecture with both convolutional and recurrent neural networks. It leverages pre-trained BERT embeddings alongside visual features, enabling multimodal tasks such as visual question answering and image captioning.
@@
VilBERT: Vision and Language BERT (VilBERT) (Lu et al., 2019) model is a pioneering architecture that integrates visual and textual information within a single BERT-based framework. Through pre-training on vast amounts of multimodal data, ViLBERT captures nuanced relationships between vision and language, enabling it to excel in tasks like visual question answering, image-text retrieval, and image captioning.
@@
VilT: The VilT (Vision-and-Language Transformer) (Kim et al., 2021) model harnesses the power of transformer architecture to seamlessly fuse textual and visual information. By jointly pre-trained on different vision and language data, VilT achieves remarkable capabilities in tasks like image captioning, visual question answering, and cross-modal retrieval, offering a versatile and powerful solution for bridging the gap between vision and language domains.
@@
SCBD: SCBD (SSE-Cross-BERT-DenseNet) (Abavisani et al., 2020) model presents an approach to real-time event detection, particularly in emergency response scenarios. The model introduces a cross-attention module to filter out irrelevant components from text and image inputs, addressing the limitations of unimodal approaches. Additionally, SCBD employs a multimodal graph-based approach, allowing stochastic transitions between embeddings of different multimodal pairs during training to improve regularization and handle limited training data effectively.
@@
MANN: Multimodal Adversarial Neural Network (MANN) (Gao et al., 2020) addresses a significant challenge in disaster-related message detection on social media by introducing an approach to identifying posts related to emerging disaster events. It has the ability to extract comprehensive features from both text and images, enabling accurate detection of disaster-related content and adaptability to unforeseen events through adversarial training.
@@
CAMM: The Cross-Attention Multi-Modal (CAMM) (Khattar and Quadri, 2022) deep neural network is proposed for classifying multimodal disaster data on social media. Integrating text and image information from Twitter posts using attention mechanisms, CAMM employs the attention mask from the textual modality to highlight relevant features in the visual modality. This cross-attention-based method emphasizes its efficacy in capturing pertinent cross-domain features in both text and image data.
@@
RoBERTaMFT: RoBERTa-based multimodal fusion transformer (Yoon et al., 2023), addresses the challenges of multimodal learning. Its co-learning methodology effectively addresses issues related to recognition performance limitations and data imbalances among modalities. RoBERTaMFT is poised to offer significant advancements in multimodal understanding and classification by promoting balanced and effective learning across different data modalities.
DMCC: Deep Multimodal Crisis Categorization Framework (DMCC) (Rezk et al., 2023) employs a two-level fusion strategy to seamlessly integrate textual and visual data. At the feature level, the MCA block ensures an effective fusion of these modalities. This holds the potential to enhance real-time crisis assessment and resource allocation, offering valuable support to emergency responders in dynamic and evolving situations.
@@ 5.5.1 Experimental Setup
In this study, all neural models are implemented using the PyTorch and TensorFlow frameworks, executed on a Linux server equipped with an Intel(R) Xeon(R) Silver 4215R CPU @3.20 GHz, 256GB of RAM, and a 16GB NVIDIA Tesla T4 GPU. The pre-trained BERT, CLIP, and ResNet models are accessible through the Hugging Face repository. We adopt the AdamW optimizer and set the learning rate hyper-parameter for the model to 5e55e-5. For the Feature Extraction and Interaction module (FEIM), we configure the dimensionality of hidden representations to $d=768$ for the textual modality and $1024$ for the visual modality. Similarly, for the Multimodal Graph Learning module (MGLM), the dimensionality of hidden representations for both textual and visual modalities is set to $d=512$. To mitigate overfitting, a dropout rate of $0.2$ is applied. The training process is conducted over 100 epochs.
@@
We systematically explore hyperparameters for our proposed model, such as learning rate, weight decay, epochs, dropout, batch size, and optimizer, with the goal of identifying optimal configurations for improved training and overall performance. We conduct comprehensive experiments to determine the most effective hyperparameter values. For the learning rate and weight decay, we test a range from 5e35e-3 to 5e75e-7, with 5e55e-5 identified as optimal. The number of epochs vary from $50$ to $120$, with $100$ epochs selected for optimal model convergence and performance enhancement. Batch sizes of $16,32,64$, and $128$ are evaluated, with a batch size of $32$ yielding the best results, balancing computational resources and memory constraints effectively. Exploration of dropout rates from $0.1$ to $0.5$ reveals $0.2$ as optimal, providing effective regularization without compromising learning capacity. Finally, after evaluating various optimizer options such as Adam, AdamW, and RMSProp, AdamW was chosen for its significant performance in optimizing the learning process. After rigorous experimentation, we determine the optimal hyperparameter configuration for our model, comprising a learning rate of 5e55e-5, batch size of $32$, $100$ epochs, a dropout rate of $0.2$, and utilization of the AdamW optimizer. These selections aim to achieve the best balance in model performance.
The CrisisSpot network comprises several key components, each with its unique set of network parameters. Firstly, the Feature Extraction and Interaction Module (FEIM) employs one linear layer for projecting the textual features in the shared embedding space for the effective alignment of the dimensions as that of visual modality. We then employ 2 dense layers with Rectified Linear Unit (ReLU) activation functions for nonlinear transformations with output dimensions of 1024 each. Each layer is followed by a dropout layer of 0.2. The final concatenated vector, Fused Attended Vector (FAV), with dimensions of 2048, is passed to the multimodal adaptive network where it initially undergoes a hierarchical reduction in dimensionality. It passes through successive fully connected layers, achieving dimensions of 1024, 512, and 256, respectively. Each layer incorporates a dropout rate of 0.2. The refined output is then directed to the Joint Fusion Learning Network (JFLN) for further processing. The FEIM component plays a crucial role in extracting and interacting with features pertinent to crisis analysis.
Next, the Multimodal Graph Learning Module (MGLM) consists of two graph neural network layers, one for each modality. The output dimension of the graph neural network layers is set to 512. These layers are followed by a dropout rate of 0.2 after each layer. Following the concatenation process, the unified vector achieves a dimensionality of 1024. Subsequently, we utilize two consecutive fully connected layers: the first layer reduces the dimensionality to 512 with a dropout rate of 0.2, followed by a second layer, further reducing it to 256 with the same dropout rate. This processed vector is subsequently directed to the JFLN, enabling the seamless integration of information from diverse sources.
The Social Context Feature Module (SCFM) utilizes two dense layers with a ReLU activation function to capture social context features. The output dimension of the first layer is set to 21, the second layer reduces the dimensions to 16. These features incorporate relevant social context information into the predictions.
Lastly, the Multimodal Fusion Network Module (MFNM) integrates features from all the modules. We employ four dense layers with output dimensions of 528, 256, 128, and 64, respectively maintaining the dropout rate of 0.2 after each layer. The final layer employs a softmax activation function, with its output dimensions dependent on the number of classes being predicted. These selected network parameters ensure that the CrisisSpot network can effectively extract, process, and fuse information from multiple sources, enabling accurate crisis analysis and prediction.
@@ 5.5.2 Performance Comparison on CrisisMMD
To affirm the effectiveness of CrisisSpot in addressing the Informative and Humanitarian Tasks on the CrisisMMD dataset (Alam et al., 2018b), we carry out a thorough comparative analysis against unimodal and multimodal state-of-the-art approaches. We measure the performance of these methods using well-established metrics as discussed in subsubsection 5.5.2. We concentrate our comparison on methods adhering to the original data split provided by the CrisisNLP repository. The results present in Table 6 demonstrate that CrisisSpot significantly outperforms both the unimodal and multimodal state-of-the-art methods. Our proposed approach yields impressive results, with an accuracy of 97.58% and an F1-score of 98.23% for the Informative Task.
Table 6: Performance Comparison on CrisisMMD for Informative and Humanitarian Tasks
Method Informative Task Humanitarian Task
Acc Prec Rec F1 Acc Prec Rec F1
BERT (Devlin et al., 2019) 86.22 86.01 86.22 86.11 81.25 82.13 82.72 82.42
ResNet (He et al., 2016) 84.88 84.76 84.89 84.82 81.33 81.29 81.98 81.63
Visual BERT (Li et al., 2019) 88.10 86.29 87.23 86.76 84.05 82.22 83.56 82.88
VilT (Kim et al., 2021) 87.60 88.29 89.18 88.73 83.30 83.82 84.92 84.37
ViLBERT (Lu et al., 2019) 88.25 88.92 90.14 89.53 84.33 84.95 86.52 85.73
SCBD (Abavisani et al., 2020) 88.52 91.13 91.84 91.48 80.62 80.77 80.52 80.64
MANN (Gao et al., 2020) 84.09 86.38 90.58 88.43 78.42 78.59 78.42 78.50
CAMM (Khattar and Quadri, 2022) 87.22 89.56 91.65 90.59 82.19 82.31 81.88 82.09
@@
RoBERTaMFT (Yoon et al., 2023) 88.90 - - 87.40 84.90 - - 80.40
DMCC (Rezk et al., 2023) 92.24 92.24 92.23 92.23 88.00 87.95 87.80 87.72
CrisisSpot 97.58 96.70 99.80 98.23 90.01 90.87 90.04 90.13
As depicted in Table 6, the performance of pre-trained unimodal methods, such as, BERT and ResNet is subpar as they are designed to process either text or image data, thus lacking the capability to harness multimodal information. VisualBERT (Li et al., 2019) underperforms due to its reliance on a single self-attention mechanism for handling both text and image data, potentially hindering its ability to capture intricate relationships between modalities. VilT (Kim et al., 2021) faces challenges due to the absence of convolutional layers and reliance solely on transformers, impacting its ability to capture local features essential in computer vision tasks. ViLBERT inherits positional encoding from BERT, potentially limiting its effectiveness in capturing spatial relationships within images compared to convolutional layers. Additionally, ViLBERT’s limited attention for cross-modality interactions may hinder its performance. SCBD (Abavisani et al., 2020) utilizes SSE-Graph, which relies on the knowledge graph for swapping embeddings. However, it may not explicitly capture the complex structural relationships within the graph. Additionally, depending on the size of the knowledge graph and the frequency of swapping, SSE-Graph may face scalability challenges. Handling large-scale graphs with frequent embedding swaps might lead to increased computational complexity. CAMM (Khattar and Quadri, 2022) uses Bi-LSTMs to process text and VGG-16 for image feature extraction. It requires access to the entire input sequence before making predictions, which can be impractical for real-time tasks requiring low latency. MANN (Gao et al., 2020) hinges on the quality of adversarial examples generated. If these examples aren’t well-crafted representations of real-world challenges, the model might not learn robust features. DMCC (Rezk et al., 2023) framework relies on the text modality’s attention mask to highlight relevant areas in the image. This may lead to errors if the text doesn’t perfectly correspond to the image content.
We attempt to incorporate various techniques to address the limitations of state-of-the-art methods. We attribute CrisisSpot’s performance enhancement to the incorporation of graph-based learning, Social Context Features, and Inverted Dual Embedded Attention which enriches the feature representation of modalities. Our proposed attention mechanism, IDEA, enables CrisisSpot to capture both similar and dissimilar pairs more effectively, facilitating its ability to attend to essential patterns. Our proposed method captures both modality-specific features and cross-modal relationships, contributing to the improved performance of our model. Furthermore, the integration of Social Contextual Features enhance the model’s context understanding by augmenting it with text-centric and user-centric scores. The various features such as User Informative Score (UIS), Crisis Informative Score (CIS), emotional context, and user engagement history enhance the classification outcomes. UIS gauges the proportion of informative content posted by a user compared to their total volume of posted content. A UIS score of 1 denotes a user has the polarity towards posting highly informative content, while 0 indicates the user’s tendency to post non-informative content. This score is integrated into the classification model as a feature, adjusting the weight of each user’s contributions during classification. Also, the Crisis Informative Score (CIS) aids in classification by quantifying the semantic attributes of the text, thereby facilitating the identification of crisis-related vocabulary within the text. Additionally, we leverage data on user engagement history to assess their influence. This includes factors such as friend count, reply count, number of likes, and other user features to further understand individual engagement patterns.
Also, it has been observed that emotional context significantly shapes classification outcomes. Through emotion and sentiment analysis techniques, we discern the emotional tone of content by analyzing sentiment polarity and identifying specific emotions expressed, including joy, sadness, anger, or fear. By incorporating sentiment and emotional cues, our classification outcomes are refined, enabling a more nuanced interpretation of user-generated content. The integration of these features forms the foundation of a robust classification framework.
5.5.3 Performance Comparison of TSEqD
To evaluate the effectiveness of CrisisSpot, we conduct a comparative analysis against existing unimodal and multimodal state-of-the-art benchmarks using our TSEqD dataset. CrisisSpot’s performance is assessed against a variety of methods, both unimodal and multimodal, to comprehensively gauge its efficacy in disaster content classification. Despite their widespread use, unimodal methods, such as, BERT and ResNet suffer from inherent limitations. BERT, originally designed for textual data, lacks the ability to process visual information, while ResNet, a convolutional neural network (CNN), operates solely on images, overlooking textual cues crucial for disaster content understanding. Multimodal methods such as ViLBERT, VisualBERT, and VILT introduce their own set of challenges. ViLBERT’s attention mechanism may not adequately handle cross-modal interactions, impacting performance. VisualBERT, though promising in its integration of both textual and visual information, faces limitations due to its reliance on a single self-attention mechanism for processing both modalities, potentially leading to suboptimal feature extraction. VILT, exclusively reliant on transformer architectures, lacks convolutional layers essential for capturing local image features, potentially limiting its ability to discern fine-grained visual details.
Table 7: Performance Comparison on TSEqD for Informative and Humanitarian Tasks
Method Informative Task Humanitarian Task
Acc Prec Rec F1 Acc Prec Rec F1
BERT (Devlin et al., 2019) 76.42 78.29 79.35 78.81 72.22 74.44 75.84 75.13
ResNet (He et al., 2016) 77.85 79.54 80.23 79.89 72.85 75.54 76.23 75.88
Visual BERT (Li et al., 2019) 78.56 80.21 81.28 80.74 74.56 76.21 78.28 77.23
VilT (Kim et al., 2021) 78.92 79.65 81.28 80.45 74.02 75.23 78.25 76.71
ViLBERT (Lu et al., 2019) 79.99 82.24 86.86 84.48 76.00 78.27 80.19 79.22
SCBD (Abavisani et al., 2020) 80.79 83.16 89.58 86.25 77.50 78.03 78.58 78.30
MANN (Gao et al., 2020) 80.14 82.76 88.55 85.55 76.80 77.25 78.95 78.09
CAMM (Khattar and Quadri, 2022) 81.08 83.89 90.34 86.99 78.00 79.43 80.34 79.88